OpenStudy (anonymous):

please help? find S8 for the geometric series 3 + -6 + 12 + -24 +…

OpenStudy (anonymous):

I know r=-2 but I don't get how to solve it

OpenStudy (campbell_st):

ok... so the sum of a geometric series is \[S_{n} = \frac{a(1 - r^n)}{(1 - r)}\] you know r = -2, n = 8 and a = 3 just substitute them all and evaluate

OpenStudy (anonymous):

I keep getting -256 and that's not an option, -255 is.. am I doing it wrong?

OpenStudy (campbell_st):

wow... I just got -255 at the 1st try

OpenStudy (campbell_st):

its \[S_{8} = \frac{3(1 - (-2)^8)}{1 - (-2)}\]

OpenStudy (anonymous):

ohh ok I got it thank you